# Context Matters: Learning Generalizable Rewards via Calibrated Features

### v2

@@ C.1.2 Equivalence threshold selection
We observe nearly identical trends across the three tested thresholds, including at key points where the learned representations are used for downstream tasks (i.e., at 100 queries per calibrated feature and 300 queries for each multi-task baseline). Based on this, we select $\epsilon=0.01$, which is the coarsest of the three thresholds and provides the most realistic approximation of human equivalence judgments.
C.1.3 Ground truth calibrated feature function descriptions
Ground truth calibrated feature functions were defined to represent complex three-dimensional relationships governing how context input should shape base features. Intuitively, picking a static context represents taking a slice of through a defined calibrated feature function, demonstrating how the calibrated feature varies with the input base feature under that context.
We define ground truth calibrated features such that they are between 0 and 1 under expected input conditions, when contextual element $c$ and input feature value $\phi$ are in their expected ranges, $c\in[0,1],\phi\in[0,1]$. Importantly, we define ground truth calibrated feature functions to describe how a single contextual element of the state affects a base feature, however our method can handle any subset of the state contextually affecting a base feature, not just a single element. We do so in this work for simplicity of illustration.
We visualize the ground truth calibrated features as a function of the base feature and affecting contextual element in Fig. 10. The actual functions being learned by calibrated feature networks are more complex, given they must not only learn the visualized relationships, but additionally invariance to all other non-relevant dimensions of the state. Descpriptions of the calibrated features for each environment are as follows (with stove distance as affected by stove heat found in all three environments):
  • All environments include the stove distance feature, which is contextually affected by stove heat. The stove distance is evenly respected across all states when the stove is off, however at only the slightest increase in temperature, only a wide berth best respects stove distance.
  • Weighted block environment includes the impact of block weight on both the table distance and the laptop distance. These relationships have similar shapes, though with different implications. For table distance, being far away (larger distance values) violates the feature more when the block is heavy. Once within a certain threshold of the table, any weight equally respects the feature value, and for light blocks, any distance is fine. For laptop distance, being close to directly above the laptop is acceptable for lighter objects, while heavier objects require a wider berth.
  • Cup environment describes the effects of cup fullness on both the cup angle and the laptop distance. Both feature functions are defined very similarly and conservatively—in most scenarios where the cup is at all full, only the widest berth fully respects the feature. The laptop distance is slightly less conservative, demonstrating that for less full cups, there is a slightly larger acceptable berth that fully respects the laptop distance constraint. Only when the cup is empty are both features equally respected everywhere.
  • Utensil environment showcases the result of utensil sharpness on both the point at the human and the distance to the human. As defined, the utensil point calibrated feature deters the utensil from ever being fully pointed at the human, even when dull. There is a sharp cutoff of about halfway, when the utensil points at 90 degrees to the person, when the feature is fully respected. When the utensil is pointing slightly at the person, the feature is more respected when the utensil is dull. The human distance calibrated feature demonstrates a preference for a barrier to be maintained away from the person, regardless of he utensil. This barrier is closer if the utensil is duller, and once past a certain sharpness this barrier is all the way at the edge of the reachable space.
@@ C.1.3 Ground truth calibrated feature function descriptions
Figure 10: Ground truth calibrated feature functions with respect to base feature and affecting contextual element.
C.1.4 Ground truth calibrated feature function definitions
@@
To help define calibrated feature functions, we first define a handful of base functions that are used to compose the ground truth calibrated feature functions:
  1. 1.
@@
    $G(\phi,c;\sigma,\mu_{\phi},\mu_{c},N,b)=\frac{N}{2\pi\sigma^{2}}\exp\left(-\frac{(\phi-\mu_{\phi})^{2}+(c-\mu_{c})^{2}}{2\sigma^{2}}\right)+b,$ (17)
@@
All composed functions follow the general form:
@@
    $f_{\text{stove}}(\phi,c)=\begin{cases}1&\text{if}\ c=0\\ \min(\max(B(\phi,c;1,2,-1,1,-3),0),1)&\text{if}\ c>0\end{cases}$ (21)
  2. 2.
@@
    $f_{\text{laptop (weight)}}(\phi,c)=\begin{cases}1&\text{if}\ c=0\ \text{or}\ \phi\geq 1\\ \min(\max(B(\phi,c;2.5,2.5,0,1,-1),0),1)&\text{otherwise}\end{cases}$ (23)
  4. 4.
@@
    $f_{\text{cup angle}}(\phi,c)=\begin{cases}1&\text{if }c=0\ \text{or}\ \phi\geq 1\\ \min\left(\max\left(\begin{aligned} &X(\phi;-2,-1.1,2,-0.65)\\ &\quad+B(\phi,c;0.5,1.2,-0.2,1,-1)\end{aligned},0\right),1\right)&\text{otherwise}\end{cases}$ (24)
  5. 5.
@@
    $f_{\text{laptop (fullness)}}(\phi,c)=\begin{cases}1&\text{if}\ c=0\ \text{or}\ \phi\geq 1\\ \min(\max(B(\phi,c;2,1.5,0,1,-1.5),0),1)&\text{otherwise}\end{cases}$ (25)
  6. 6.
@@
    $f_{\text{human}}(\phi,c)=\begin{cases}1&\text{if }\phi\geq 1\\ \min\left(\max\left(\begin{aligned} &G(\phi,c;0.2,1,0,6,-0.55)\\ &\quad+X(\phi;-2,-1.1,2,-0.4)\end{aligned},0\right),1\right)&\text{otherwise}\end{cases}$ (26)
  7. 7.
@@
    $f_{\text{point}}(\phi,c)=\begin{cases}1&\text{if }\phi<\frac{1}{3}\\ \min\left(\max\left(\begin{aligned} &X(\phi;0.2,-1,0.5,0)\\ &\quad+G(\phi,c;0.5,0.4,0,2,-0.5)\end{aligned},0\right),1\right)&\text{otherwise}\end{cases}$ (27)
C.2 Training and evaluation
C.2.1 Training
We use a two-stage training process: first learning intermediate representations from training queries, and then training reward functions on top of those representations. For each of six random seeds, we independently split the environment state set of 10k states into 80% training and 20% test states, resulting in different train/test partitions per seed. For each seed, training queries are used for both representation and reward learning, while test queries are used for evaluation only. During representation and reward training, queries are generated by randomly sampling state pairs from the training states.
C.2.2 Metrics
We use two primary metrics to evaluate representations and rewards: calibrated feature accuracy (MSE) and reward accuracy. We evaluate calibrated feature accuracy by computing the mean squared error (MSE) between predicted and ground-truth feature values across the 2,000 test states per seed. Predicted calibrated feature values are normalized between 0 and 1 based on the min and max logits observed during training. Occasionally, especially with untrained models (e.g., 0 training queries), predictions fall outside the logit bounds seen during training (default min and max bounds of 0 and 1), leading to MSE values greater than 1 (see examples in Fig. 12). To evaluate reward functions, we sample 1,000 state pairs from the test set for each seed. We compute the percentage of correctly predicted pairwise preferences according to the ground-truth reward function. This metric is used for both direct evaluation of reward heads trained during representation learning and for downstream reward functions trained on top of frozen or fine-tuned representations. If the ground-truth rewards of a test comparison state pair are within the selected equivalence threshold (App. C.1.2), that pair is excluded from evaluation. Across all 720 test comparison state sets (comprising 10 test rewards evaluated for the full calibrated features reward and each of the 3 single-feature calibrated rewards, across 3 environments and 6 seeds), the number of non-equivalent comparison pairs used for evaluation (out of 1000 possible) had a median of 1000, mean of 956 (SD = 108), and a minimum of 394. Only about 5% of test sets had fewer than 700 usable comparison pairs, with over 80% of those cases stemming from the single test reward in which only one feature was relevant.
C.2.3 Evaluation
Calibrated feature representation learning. For representations that output calibrated features, we evaluate the learned features using the MSE metric described above. This allows us to assess how accurately each feature dimension captures its intended ground-truth signal, even with varying amounts of training data.
Multi-Task representation learning. For multi-task baselines, we evaluate reward accuracy for each of the $N$ training rewards on the same 1,000 test state pairs (excluding equivalent pairs under the training reward). We then average the reward accuracy across the $N$ heads for each seed. This measures how well the shared representation supports multiple reward functions simultaneously.
Downstream reward learning. To assess generalization, we train 10 new (held-out) reward functions on top of the learned representations. These rewards are trained either with the learned representation frozen or unfrozen. Each reward is evaluated on the same 1,000 test state pairs (excluding equivalent pairs), and we report the average reward accuracy across the 10 test rewards for each seed.
C.2.4 Aggregation and error bars
All results are averaged over six seeds. For reward accuracy, we first average across the evaluated rewards ($N$ multi-task training rewards or 10 downstream test rewards) within each seed, then report the mean and standard error across seeds. Therefore, error bars shown in plots reflect standard error across the seed variation, not reward variation.
C.3 Representation learning results
We first present representation learning results for both calibrated feature models and multi-task baselines. For calibrated features, we plot the mean squared error (MSE) of all three learned calibrated features for each environment as a function of the number of training queries (Fig. 11). Each curve represents the MSE for one calibrated feature, evaluated on held-out test states. We also include a zoomed-out version of these plots (Fig. 12), where a dashed horizontal line indicates the MSE of a randomly initialized model (computed as the average MSE across the three features at zero training queries). Calibrated feature MSE can struggle to approach zero when learning from comparisons, which only provide ordinal relationships between pairs of points rather than absolute function values. Some smoother functions (such as cup and laptop distance, which have similar defined calibrated feature functions) are easier to capture accurately with only pairwise comparisons, while others (e.g., distance to human) can struggle to achieve low MSE (Fig. 11).
We next present multi-task representation learning results, which report reward accuracy averaged over the $N$ training rewards in each environment (Fig. 1314). Notably, for multi-task baselines, separate representation learning was required for each downstream reward combination – for each single calibrated feature scenario and for the all calibrated features scenario – since the shared representation depends on the specific reward training setup. In contrast, our calibrated feature approach supports reuse: the same three calibrated features can be applied across all downstream reward configurations in an environment, enabling flexible generalization without retraining the representation. We can see that as expected, the SinglePref model achieves high reward accuracy the most rapidly, given it is only trained and evaluated on a single reward. In contrast, as the number of $N$ training tasks for each JointPref method increases, the more queries it takes to find a shared representation that supports accurate rewards across all training tasks.
@@ C.3 Representation learning results
We select 100 pre-training queries given we see general convergence behavior for calibrated features at this point. When all three features are contextually-affected we use 300 pre-training queries so that all calibrated features still see 100 pre-training queries when divided among the three features.
Figure 11: Calibrated feature learning results.
Figure 12: Fig. 11 results compared to random network initialization.
Figure 13: Multi-task representation learning results for multiple contextually-affected feature experiments.
Figure 14: Multi-task representation learning results for single contextually-affected feature experiments.
C.4 Reward learning with and without frozen representations
@@
We visualize downstream reward learning results (Fig. 1516), displaying performance when intermediate representations are either frozen or fine-tuned during reward training. For our method, we observe that freezing the learned calibrated feature representations consistently yields higher reward accuracy, particularly in low-data regimes. This suggests that our method successfully learns meaningful and reusable intermediate features that generalize well to new tasks. In contrast, for the multi-task baseline, freezing and unfreezing perform similarly at low query counts, with frozen representations sometimes performing slightly better. However, as more data becomes available, fine-tuning the representation eventually leads to higher performance, indicating that the multi-task representation benefits from task-specific adaptation. In the main text we show the frozen variant of our method and the unfrozen variants of all multi-task baseline methods.
To quantify the overall performance gap, we compute the difference between the best-performing version of our method (always frozen) and the best-performing multi-task baseline configuration (which varies by $N$ training rewards and freezing condition). We summarize these comparisons in Table 3, reporting the also the relative increase in reward accuracy (increase in percentage of correct query selection) demonstrated by our method at both 5 and 10 reward queries. In these low-data regimes, our method achieves improvements of up to 14.91% in reward accuracy over the highest-performing baseline.
Env. Calib. Feats CF 5q JointPref 5q $\Delta\%$ 5q CF 10q JointPref 10q $\Delta\%$ 10q
Weighted All 0.808 0.742 +6.6 0.848 0.773 +7.5
Block Stove 0.802 0.653 +14.9 0.841 0.698 +14.3
Table 0.804 0.663 +14.1 0.844 0.697 +14.7
Laptop 0.794 0.665 +12.9 0.828 0.705 +12.3
Cup All 0.807 0.714 +9.3 0.824 0.726 +9.8
Stove 0.747 0.653 +9.4 0.813 0.692 +12.1
Cup Angle 0.763 0.660 +10.3 0.835 0.688 +14.7
Laptop 0.814 0.690 +12.4 0.840 0.731 +10.9
Utensil All 0.793 0.754 +3.9 0.823 0.775 +4.8
Stove 0.855 0.733 +12.2 0.867 0.783 +8.4
Human 0.827 0.691 +13.6 0.853 0.762 +9.1
Point 0.802 0.743 +5.9 0.825 0.772 +5.3
Table 3: Reward accuracy comparison between calibrated features and the best-performing multi-task variant at 5 and 10 training queries. $\Delta\%$ indicates percent improvement in reward accuracy.
C.5 Qualitative visualization details
The qualitative visualizations display learned calibrated feature values at 8,000 different EE locations across four distinct contexts. The visualization is limited to 8,000 points due to PyBullet’s rendering constraints. Visualization states are generated similar to how described before in App. B, by sampling random points within the robot’s reachable workspace and solving inverse kinematics to obtain joint angles (with object positions remaining constant). However, rather than using random contexts, each point is evaluated across the discrete contexts defined at the end of App. B. For each visualized feature, the relevant contextual element is systematically varied across its defined discrete values while the calibrated feature is evaluated for all points in the set. Non-relevant contextual elements are held constant at zero.
Feature values are depicted as points at each EE location, where lighter colors indicate higher feature values. Learned calibrated feature values are normalized across the entire point cloud evaluated over all discretized contextual values, with non-affecting contextual features held constant. We then select four spread out contexts to visualize the point cloud evaluation in our final figures. All visualizations use calibrated features trained with seed=0 for consistency. We visualize just four features given these visualizations work well for 3D distance-based relationships, but become more challenging for complex transformations like rotations, which would benefit from projection onto relevant 2D planes.
Figure 15: Multiple contextually-affected feature experiments shown with frozen and unfrozen representations.
@@ C.5 Qualitative visualization details
Figure 16: Single contextually-affected feature experiments shown with frozen and unfrozen representations.
@@ Appendix D User Study
D.1 Result details
The visualized box and whisker plots in 5 display the distribution of data through five key summary statistics. The box represents the interquartile range (IQR) with the bottom edge at the 25th percentile and top edge at the 75th percentile, while a line inside the box marks the median (50th percentile). The whiskers extend from the box edges to the most extreme data points that fall within 1.5 times the IQR from the box boundaries. Any data points beyond this range are plotted as individual outlier points, shown as small circles.
Further user study statistics can be visualized in Table 4. Statistics are shown for both test features individually, as well as across both features. Items on the 7-point Likert scale were summed to represent a composite score representing general contextual preference alignment, making the possible Likert scale score range 4 to 28, given 4 Likert items. Ranking scores range from 1 to 4, with 1 being the highest, or best model, and 4 being the lowest. We can see that rankings and Likert scores averaged across both features increase with the number of training queries used for a model, with a leveling off between 50 and 100 queries. The same trend is observed when individually inspecting the human distance feature. The stove distance feature shows evidence of flatter leveling off, with models trained with 50 and 100 queries having equivalent average Likert scale scores, and the model with 50 queries having a slightly higher average ranking than 100 queries.
Table 4: Mean and sample standard deviation of Likert and Ranking scores for calibrated features trained on different query counts.
Feature Metric 0 Queries 25 Queries 50 Queries 100 Queries
Stove Distance Likert Score 13.67 $\pm$ 6.92 18.25 $\pm$ 5.61 22.58 $\pm$ 3.63 22.58 $\pm$ 4.01
Ranking 3.08 $\pm$ 1.24 3.08 $\pm$ 0.79 1.83 $\pm$ 0.94 2.00 $\pm$ 0.95
Human Distance Likert Score 10.92 $\pm$ 5.57 20.50 $\pm$ 3.45 22.00 $\pm$ 4.65 22.83 $\pm$ 5.54
Ranking 3.50 $\pm$ 1.00 2.75 $\pm$ 0.87 2.17 $\pm$ 1.03 1.58 $\pm$ 0.67
Both Features Likert Score 12.29 $\pm$ 6.30 19.38 $\pm$ 4.70 22.29 $\pm$ 4.09 22.71 $\pm$ 4.73
Ranking 3.29 $\pm$ 1.12 2.92 $\pm$ 0.83 2.0 $\pm$ 0.98 1.79 $\pm$ 0.83
D.2 Preference analysis
Each participant provided a short free-response description of their contextual preference prior to responding to contextual feature queries (see prompt in Fig. 20). Therefore, we collected 12 short free-response answers regarding contextual preferences for each of the two test features. A single reviewer manually reviewed the responses and conservatively grouped them based on the level of specificity provided (Tables 56). Responses describing contextual preferences in general, high-level terms without specific details were grouped together, while responses including concrete details were grouped separately unless the details matched exactly. This inductive grouping was intended to get a preliminary understanding of preference diversity. We found 6 users shared a general preference for the human distance feature, with 2 users sharing a similar variant and 4 users holding unique variants. For the stove distance feature, we found 7 users shared a general preference, and the remaining 5 users described individual variations.
Figure 17: Differences in user preferences captured in learned calibrated features.
In Fig 17 we give another example similar to Fig 6, where we demonstrate how learned calibrated features differ for users with different preferences. The main difference in this example is the context in which the stove is off – while for one user the feature is equally respected in all states when the stove is off, the other user still desires the robot to stay away from directly above the stove, even when it is off.
D.3 Environments
The same environments are used, specifically the Cup environment was used for the training feature and the Utensil environment was used for the two test features, however with discrete contexts for visualization purposes as defined at the end of App.B. Environment state sets of 10k states (with discrete contexts) were generated, and a set of 100 queries (50 for training feature) were randomly sampled and maintained consistent across participants. The same state pairs were shown for training both test features in the utensil environment. Additionally, while the laptop location was represented in the state, when training the two test features in the Utensil environment, the irrelevant laptop was not visualized in order to prevent cluttering of the visualization.
Figure 18: Query response GUI. State 1 shown for first state pair in comparison set, which was the same across both experiment features. Participants visualized and responded to queries using buttons on right panel. Left hand panel is irrelevant and can be ignored.
Figure 19: Model visualization GUI. Example shown of calibrated feature trained for stove distance with 50 contextual feature queries. Visualized in high heat context. Participants inspected different models using buttons on right panel. Left hand panel is irrelevant and can be ignored.
D.4 Graphical user interfaces
Participants interacted with two graphical user interfaces (GUIs) in order to both respond to contextual feature queries (Fig. 18), and to interactively inspect learned calibrated features (Fig. 19). When responding to queries, participants were able to view the states from different angles, and in response to the visualized contextual feature query select between "Select State 1", "Select State 2", or "States are equal". Users could also switch back and forth between the two compared states, and the current state was visualized in order to help users stay aware. Additionally, environment objects, like utensils, were colored differently to help make relevant context immediately salient and speed up query response.
When visualizing and interactively inspecting calibrated features, participants were presented with four models ("Relationships" 1-4) which represented calibrated features trained on 0 (base feature), 25, 50, and all 100 labeled queries. Learned models were visualized just as described in App. C.5, however 5000 states were used as opposed to 8000 for speed of evaluation and visualization. Participants could switch between the four models, and visualize a model in varying contexts by dragging a sliding bar which snapped the visualized context to the nearest of four spread out discrete contexts (and changed the model visualization accordingly). Participants were instructed on helpful ways to investigate and compare models, such as inspecting a single models across contexts, and switching between all models to see how they compare in a single context. Participants provided no input in this GUI, instead using it to respond to a survey collecting the Likert scale evaluation for each model and model rankings.
D.5 Procedures
Participants completed a familiarization phase, where they were trained in contextual feature queries and practiced providing 50 contextual feature queries for the laptop distance feature in the Cup environment. Participants also practiced visualizing and inspecting three learned models trained on 0 (base feature), 25, and all 50 of their provided practice queries. Participants then completed two experiment rounds of query response, which only differed from the practice round in the number of queries answered (100), and the resulting number of models visualized (4 instead of 3).

### v3

@@ C.1.1. Equivalence threshold selection
We observe nearly identical trends across the three tested thresholds, including at key points where the learned representations are used for downstream tasks (i.e., at 100 queries per calibrated feature and 100/300 queries for each multi-task baseline). Based on this, we select $\epsilon=0.01$, which is the coarsest of the three thresholds and provides the most realistic approximation of human equivalence judgments.
Line plots showing calibrated feature MSE against number of representation learning queries. Three lines shown for three different equivalence thresholds (0.01, 0.001, 0.0001). Line plots show minimal difference between the three lines.
Figure 8. Equivalence threshold experiments for calibrated features.
Line plots showing calibrated feature MSE against number of representation learning queries. Three lines shown for three different equivalence thresholds (0.01, 0.001, 0.0001). Line plots show minimal difference between the three lines.
Line plots showing multi-task training rewards against number of representation learning queries. Three lines shown in each plot for three different equivalence thresholds (0.01, 0.001, 0.0001). One plot shown for each multi-task variant. Line plots show minimal difference between the three lines in each plot.
Figure 9. Equivalence threshold experiments for multi-task baselines.
Line plots showing multi-task training rewards against number of representation learning queries. Three lines shown in each plot for three different equivalence thresholds (0.01, 0.001, 0.0001). One plot shown for each multi-task variant. Line plots show minimal difference between the three lines in each plot.
3D figures showing how calibrated feature value is affected by changes in base feature and relevant contextual element.
@@
Figure 10. Ground truth calibrated feature functions with respect to base feature and affecting contextual element.
3D figures showing how calibrated feature value is affected by changes in base feature and relevant contextual element.
C.1.2. Ground truth calibrated feature function definitions
Ground truth calibrated feature functions were defined to represent complex three-dimensional relationships governing how context input should shape base features. Intuitively, picking a static context represents taking a slice of through a defined calibrated feature function, demonstrating how the calibrated feature varies with the input base feature under that context. We visualize the ground truth calibrated features as a function of the base feature and affecting contextual element in Fig. 10. The actual functions being learned by calibrated feature networks are even more complex, given they must not only learn the visualized relationships, but additionally invariance to all other non-relevant dimensions of the state.
We define ground truth calibrated features such that they are between 0 and 1 under expected input conditions, when contextual element $c$ and input feature value $\phi$ are in their expected ranges, $c\in[0,1],\phi\in[0,1]$. Importantly, we define ground truth calibrated feature functions to describe how a single contextual element of the state affects a base feature; however, our method can handle any subset of the state contextually affecting a base feature, not just a single element.
@@ C.1.2. Ground truth calibrated feature function definitions
Base functions. To help define calibrated feature functions, we first define a handful of base functions that are used to compose the ground truth calibrated feature functions:
  1. (1)
@@
    (17) $G(\phi,c;\sigma,\mu_{\phi},\mu_{c},N,b)=\frac{N}{2\pi\sigma^{2}}\exp\left(-\frac{(\phi-\mu_{\phi})^{2}+(c-\mu_{c})^{2}}{2\sigma^{2}}\right)+b,$
@@
Calibrated features. All composed functions follow the form:
@@
    (21) $f_{\text{stove}}(\phi,c)=\begin{cases}1&\text{if}\ c=0\\ \min\Bigl(\max(B(\phi,c;1,2,-1,1,-3),0),1\Bigr)&\text{if}\ c>0\end{cases}$
  2. (2)
@@
    (23) $f_{\text{laptop (weight)}}(\phi,c)=\begin{cases}1&\text{if }c=0\text{ or }\phi\geq 1\\[6.0pt] \min\Bigl(\max\bigl(\\ \hskip 5.0ptB(\phi,c;2.5,2.5,0,1,-1),0\bigr),1\Bigr)&\text{otherwise}\end{cases}$
  4. (4)
@@
    (24) $f_{\text{cup angle}}(\phi,c)=\begin{cases}1&\text{if }c=0\ \text{or}\ \phi\geq 1\\ \min\Bigl(\max\Bigl(\\ \hskip 5.0ptX(\phi;-2,-1.1,2,-0.65)\\ \hskip 5.0pt+B(\phi,c;0.5,1.2,-0.2,1,-1),0\Bigr),1\Bigr)&\text{otherwise}\end{cases}$
  5. (5)
@@
    (25) $f_{\text{laptop (fullness)}}(\phi,c)=\begin{cases}1&\text{if}\ c=0\ \text{or}\ \phi\geq 1\\ \min\Bigl(\max(\\ \hskip 5.0ptB(\phi,c;2,1.5,0,1,-1.5),0),1\Bigr)&\text{otherwise}\end{cases}$
  6. (6)
@@
    (26) $f_{\text{human}}(\phi,c)=\begin{cases}1&\text{if }\phi\geq 1\\ \min\Bigl(\max\Bigl(\\ \hskip 5.0ptG(\phi,c;0.2,1,0,6,-0.55)\\ \hskip 5.0pt+X(\phi;-2,-1.1,2,-0.4),0\Bigr),1\Bigr)&\text{otherwise}\end{cases}$
  7. (7)
@@
    (27) $f_{\text{point}}(\phi,c)=\begin{cases}1&\text{if }\phi<\frac{1}{3}\\ \min\Bigl(\max\Bigl(\\ \hskip 5.0ptX(\phi;0.2,-1,0.5,0)\\ \hskip 5.0pt+G(\phi,c;0.5,0.4,0,2,-0.5),0\Bigr),1\Bigr)&\text{otherwise.}\end{cases}$
C.2. Representation learning results
@@
We present intermediate representation learning results for our method and multi-task baselines (Fig. 1113). These visualizations were used to select the number of pre-training queries set for downstream reward learning. We select 100 pre-training queries given we see general convergence behavior for calibrated features at this point (Fig. 11). When all three features are contextually-affected we use 300 pre-training queries so that all calibrated features still see 100 pre-training queries when divided among the three features.
For calibrated features, we plot the mean squared error (MSE) of all three learned calibrated features for each environment as a function of the number of training queries (Fig. 11), evaluated across the 2k test states per seed. Calibrated feature MSE can struggle to approach zero when learning from comparisons, which only provide ordinal relationships between pairs of points rather than absolute function values. Some smoother functions (such as cup and laptop, which have similar defined calibrated feature functions) are easier to capture accurately with only pairwise comparisons, while others (e.g., human) can struggle to achieve low MSE.
We next present multi-task representation learning results, which report reward accuracy averaged over the $N$ training rewards in each environment (Fig. 1213). We can see that as expected, the SinglePref model achieves high reward accuracy the most rapidly, given it is only trained and evaluated on a single reward. In contrast, as the number of $N$ training tasks for each JointPref method increases, the more queries it takes to find a shared representation that supports accurate rewards across all training tasks. Notably, for multi-task baselines, representation learning was required to be repeated for the scenario in which all features are contextually-affected. This is because the contextual adaptations learned in each single contextually-affected feature scenario operate on the full feature set, therefore individual feature adaptations cannot be easily extracted and composed. In contrast, our calibrated feature approach supports reuse: the same three calibrated features were combined in the all calibrated features setting without retraining the representation.
Line plots showing calibrated feature MSE against number of representation learning queries. A separate plot is shown for each environment. MSE generally converges around 100 queries, except for the distance to human feature in the Utensil environment, which takes closer to 300 queries to converge.
Figure 11. Calibrated feature learning results.
Line plots showing calibrated feature MSE against number of representation learning queries. A separate plot is shown for each environment. MSE generally converges around 100 queries, except for the distance to human feature in the Utensil environment, which takes closer to 300 queries to converge.
C.3. Freezing vs. fine-tuning representations
@@
We compare reward learning when pre-trained representations are either frozen or fine-tuned (Fig. 1514). We observe that freezing the learned calibrated feature representations consistently yields higher reward accuracy, particularly in low-data regimes. This suggests that our method successfully learns meaningful intermediate features that can be reused in new rewards. In contrast, for the multi-task baseline, freezing and unfreezing perform similarly at low query counts, with frozen representations sometimes performing slightly better. However, as more data becomes available, fine-tuning the representation eventually leads to higher performance, indicating that the multi-task representation benefits from task-specific adaptation. In the main text we show the frozen variant of our method and the unfrozen variants of all multi-task baseline methods.
Line plots showing training reward accuracy against number of representation learning queries, which in this case are training preference queries. Training reward accuracy generally takes longer to converge the greater the number of training rewards there are.
Figure 12. Multi-task representation learning results for single contextually-affected feature experiments.
Line plots showing training reward accuracy against number of representation learning queries, which in this case are training preference queries. Training reward accuracy generally takes longer to converge the greater the number of training rewards there are.
Line plots showing training reward accuracy against number of representation learning queries, which in this case are training preference queries. Training reward accuracy generally takes longer to converge the greater the number of training rewards there are.
Figure 13. Multi-task representation learning results for multiple contextually-affected feature experiments.
Line plots showing training reward accuracy against number of representation learning queries, which in this case are training preference queries. Training reward accuracy generally takes longer to converge the greater the number of training rewards there are.
Line plots showing reward accuracy against number of reward preference queries. For each method a there are two lines: one with frozen representations and one with unfrozen representations. Our method performs better with frozen representations while baselines generally perform better with unfrozen representations.
@@
Figure 14. Single contextually-affected feature experiments shown with frozen and unfrozen representations.
Line plots showing reward accuracy against number of reward preference queries. For each method a there are two lines: one with frozen representations and one with unfrozen representations. Our method performs better with frozen representations while baselines generally perform better with unfrozen representations.
Line plots showing reward accuracy against number of reward preference queries. For each method a there are two lines: one with frozen representations and one with unfrozen representations. Our method performs better with frozen representations while baselines generally perform better with unfrozen representations.
Figure 15. Multiple contextually-affected feature experiments shown with frozen and unfrozen representations.
Line plots showing reward accuracy against number of reward preference queries. For each method a there are two lines: one with frozen representations and one with unfrozen representations. Our method performs better with frozen representations while baselines generally perform better with unfrozen representations.
@@ Appendix D User Study
D.1. Procedures